A leather coin purse has the shape of a triangle. Two sides are equal in length and the third side is 3cm shorter than one and one half times the length of the equal sides. the perimeter is 28.5 cm. find the lengths of the sides.
Answer by mananth(16949)

let the equal sides be x
the third side = x/2-3
..
x+x+x/2 -3=28.5
5x/2 = 28.5+3
5x/2=31.5
5x=63
/5
x=63/5
x=12.6 cm the equal sides
third side = 6.3
...
m.ananth@hotmail.ca